These are files for corrected updated version of tables of biquadratic bicyclic number fields. We give tables of class numbers and fundamental units of fields with discriminant ≦ 10^{8}. All of them are in AMS-TEX. It is easy to translate for other style of TEX, because most parts of these files are in plain style.

The files ‘rbbf8t1.tex’, ‘rbbf8t2.tex’, and ‘rbbf8t3.tex’ are files for real fields. The file ‘rbbf8t1.tex’ is the source of Table 1, which provides fundamental data for fields. The file ‘rbbf8t2.tex’ is the source of Table 2, which provides narrow class numbers and generators of the totally positive units. The file ‘rbbf8t3.tex’ is the source of Table 3, which provides explicit forms of fundamental units. The sizes of these files ammount about 1.11 MB.

The files ‘ibbf8t1.tex’and ‘ibbf8t2.tex’ are files for imaginary fields. The file ‘ibbf8t1.tex’ is the source of Table 1, which provides fundamental data for fields. The file ‘ibbf8t2.tex’ is the source of Table 3, which provides explicit forms of fundamental units. The sizes of these files ammount about 1.99 MB.

In the file ‘construct.tex’, it is described how to construct tables roughly.

The file ‘correction.tex’ is the correction to the previous version of tables.

Typeset by AMS-TEX